OpenStudy (anonymous):

solve for x and y 3x - 2y = 15 3x + 2y = 7

OpenStudy (anonymous):

eq1. 3x-2y = 15 eq2. 3x+2y = 7 (eq1 + eq2) : 6x=22 , x =22/6 (eq1 - eq2): -4y = 8, y =-2

OpenStudy (anonymous):

thank you very much